The following jobs are waiting to be processed at a small machine​ center:

JOB DUE DATE DURATION (DAYS)
10 260 8
20 252 18
30 260 26
40 270 32
50 275 14

All dates are specified as manufacturing planning calendar days. Assume that all jobs arrived on day 200 ​(yesterday) in the order shown but are not scheduled to begin until day 201 (today).

a. Sequence the jobs according to LPT?

b. What is the average completion (flow) time?

c. What is the average job lateness?

d. What is the average number of jobs in the system?

c. What is the average job lateness?

d. What is the average number of jobs in the system?
Answer #1

Answer:

Rule: The job with longest duration will pe processed first.

Flow time = Cumulative sum of duration of completed jobs

Completion Date = 200 + Flow time (As job can start on Day 201)

Lateness = Max(0, Completion Date - Due Date)

a) Sequence: 40, 30, 20, 50, 10

b) Average Flow time = Total Flow Time / 5 = 354/5 = 70.80

c) Average Job Lateness = Total Lateness/5 = 77/5 = 15.4

d) Average no. of jobs in system = Total Flow Time / Total Duration = 354/98 = 3.61

LPT Sequence Job Due Date Duration Flow time Lateness 40 0 1 2 3 30 270 260 252 275 260 32 58 76 0 Completion Date 232 258 27